dakotas math test grade was 7 points less than his science test grade. the sum of the grades was 183%. what did dakota get on his math test
Answer by rwm(914)

m=s-7
since we want the math score we solve for m
s=m+7
m+m+7=183
2m=176
m=88
btw
s=95